OpenStudy (anonymous):

Find the 50th derivative of y=cos2x ?Find the 50th derivative of y=cos2x ?@Mathematics

OpenStudy (anonymous):

\[y'=-2\sin(2x)\] \[y''=-4\cos(2x)\] \[y'''=8\sin(2x)\] \[y^{(4)}=2^4\cos(2x)\] should be a good start

myininaya (myininaya):

\[f^{(1)}(x)=-2\sin(2x)\] \[f^{(2)}(x)=-4\cos(2x)\] \[f^{(3)}(x)=8\sin(2x)\] \[f^{(4)}(x)=16\cos(2x)\] \[f^{(5)}(x)=-32\sin(2x)\] its doing this -,-,+,+,-,-,+,+,-,-,... \[f^{(50)}(x)=2^{50}\cos(2x)\] <now is this negative or positive?

myininaya (myininaya):

i will leave you to answer that question i have to go

OpenStudy (anonymous):

later!

OpenStudy (anonymous):

okay I'll try that thank you!

hero (hero):

 If you keep going counterclockwise and count to 50 you'll see that the 50th derivative of y = sin(x) is